# Lung-DDPM+: Efficient Thoracic CT Image Synthesis using Diffusion Probabilistic Model

### v1

@@ 1 Introduction
Lung cancer is one of the greatest threats to human well-being. In 2022, it remained the most frequently diagnosed cancer worldwide, with approximately 2.5 million new cases, accounting for 12.4% of all cancer diagnoses. Furthermore, lung cancer continues to be the leading cause of cancer-related mortality, responsible for approximately 1.8 million deaths (18.7% of total cancer deaths) in 2022 [1]. For small, localized tumors (stage I), surgical resection of non-small cell lung cancer (NSCLC) offers a favorable prognosis, with 5-year survival rates of up to 70%. In contrast, cases diagnosed with distant metastatic disease (stage IV) have a 1-year survival rate of only 15-19%, compared to 81-85% for stage I [2]. These statistics underscore the critical importance of early detection of lung nodules and intervention in lung cancer diagnosis.
@@
Given the rapid advancement of AI in healthcare, numerous deep learning-based models have been deployed across various domains. However, this progress exacerbates data scarcity, which significantly impacts AI-assisted diagnostic models for early lung cancer detection. Recently, generative AI has garnered increasing attention for its ability to provide high-quality synthetic samples, enhancing model performance and mitigating data limitations. In radiological imaging, generative AI applications extend to X-ray [3], computed tomography (CT) [4], positron emission tomography (PET) [5], magnetic resonance imaging (MRI) [6], and ultrasound [7]. Among these, CT image synthesis remains one of the most challenging tasks due to its high computational complexity and the intricate texture details of anatomical structures.
@@
With the emergence of generative AI, various models for CT image synthesis have been introduced over the past decade. Since 2020, our team has focused on generative models for thoracic CT synthesis, initially proposing a generative adversarial network (GAN)-based approach for 2D COVID-19 CT image synthesis [4]. More recently, we advanced this work by introducing Lung-DDPM [8], a denoising diffusion probabilistic model (DDPM) for 3D CT image synthesis with lung nodules. While Lung-DDPM addressed three key limitations of existing methods—anatomical imprecision, imperfect blending, and inconsistency—SOTA models still suffer from low efficiency in generating 3D thoracic CT images.
@@
As illustrated in Figure 1 (A), we compare the sampling times of the proposed Lung-DDPM+ with three SOTA models—Lung-DDPM [8], Med-DDPM [6], and GEM-3D [9]—for generating thoracic CT images of two different sizes. To synthesize a single CT image, SOTA generative models require anywhere from tens to over two hundred seconds, resulting in low efficiency during the sampling process. This inefficiency severely limits the practical deployment of generative models in downstream medical applications, particularly in time-sensitive and cost-sensitive scenarios.
@@
Beyond the low efficiency issue, we also observed that SOTA generative models tend to synthesize imprecise anatomical structures when no semantic layout guidance is available. In Figure 1 (B), we illustrate a case generated by Lung-DDPM and Med-DDPM, both exhibiting anatomical imprecision. Compared to the proposed Lung-DDPM+ sample, the synthetic image produced by Med-DDPM failed to maintain anatomical precision in extra-pulmonary areas. While Lung-DDPM preserved extra-pulmonary regions, its synthetic sample lacked pulmonary vessel and bronchus details. In clinical practice, experienced radiologists manually annotate detailed semantic layouts for organs and lesions. However, this process is time-consuming and costly [10], making well-labeled medical imaging datasets scarce for researchers. Therefore, developing a feasible approach to address the anatomical imprecision issue in scenarios with limited annotations remains a critical research challenge.
@@
    Our method has low data requirements and does not rely on externally sourced healthy cases, making it applicable to datasets in which all cases contain nodules (e.g., LIDC-IDRI [11]). By synthesizing additional nodule-positive samples with greater diversity, our approach enhances the robustness and performance of downstream segmentation models. While healthy cases are not required, they can be incorporated to further expand the diversity of the synthetic dataset.
@@ 2.1 Efficient denoising diffusion probabilistic models
Since the introduction of DDPM [12], it has emerged as one of the most advanced generative models due to its strong ability to generate high-quality data across various applications. However, low efficiency remains a critical limitation, restricting both its performance and deployment. To address this issue, researchers have explored multiple strategies to enhance DDPM efficiency. Improved model architectures, such as latent space diffusion models [13], allow DDPMs to operate in a compressed space, significantly reducing computational demands. Adaptive noise scheduling methods, including PriorGrad [14], introduce data-dependent priors to accelerate convergence and improve sample quality. Additionally, hybrid model integration, which combines diffusion models with pre-trained generative architectures like GANs and VAEs [15], leverages the strengths of both approaches to enhance sampling speed. Further optimizations focus on the sampling process, where techniques such as efficient sampling schedules [16] and early stopping mechanisms [17] reduce the number of denoising steps while maintaining generation quality. Despite these advancements, research on optimizing DDPM sampling efficiency in medical imaging remains limited, necessitating further exploration to improve its feasibility in clinical applications.
@@ 2.2 Generative models for medical imaging
Over the past decade, generative models have been extensively utilized in medical imaging, driven by rapid advancements in deep learning. For instance, Jiang et al. [4] employed a conditional generative adversarial network (cGAN) to synthesize COVID-19 CT images, facilitating data augmentation for improved diagnostic models. Dorjsembe et al. [6] introduced conditional diffusion models for semantic 3D brain MRI synthesis, enhancing the generation of realistic brain images for various clinical applications. Zhu et al. [9] proposed a generative enhancement approach for 3D medical images, focusing on improving image quality and resolution. Diffusion models have also gained traction in medical imaging, as highlighted in a recent review by Hein et al. [18], which explores the transformative role of physics-inspired generative models in this domain. Among these advancements, Jiang et al. [8] developed Lung-DDPM, a semantic layout-guided diffusion model for lung CT image synthesis, demonstrating superior performance over other state-of-the-art (SOTA) methods. However, like other approaches, it suffers from low efficiency, a key challenge this paper aims to address.
@@ 3.2 Pulmonary DPM-solver
Unlike Lung-DDPM, which operates on the entire pulmonary region, the proposed method focuses on the effective synthesis of nodule regions. To accelerate the anatomically aware sampling (AAS) process, we introduce a pulmonary DPM-solver, which extends the original DPM-solver [19, 20] by dynamically blending nodules with surrounding pulmonary areas during the EAAS sampling process.
@@
Despite their impressive generative capabilities, standard DPMs suffer from slow inference, often requiring hundreds or thousands of function evaluations (NFEs) for high-quality sampling. To mitigate this limitation, DPM-Solver [19] was introduced as a fast, high-order ODE solver that significantly accelerates sampling while preserving sample fidelity.
@@
Building upon DPM-Solver, DPM-Solver++ [20] further optimizes inference speed and quality by integrating both stochastic and deterministic solvers in a hybrid framework. In early diffusion steps, stochastic updates enhance sample diversity, while deterministic high-order solvers refine results in later timesteps:
@@ 3.3.1 Nodule semantic layout generation
The generation of pulmonary CT images relies heavily on high-quality lung nodule semantic layouts, which are typically created by experienced radiologists in the Lung-DDPM workflow. To reduce radiologists’ workload and enhance the efficiency of the generation process, we introduce a rule-based nodule semantic layout generation method. Specifically, following [21] and [22], we first generate realistic tumor-like shapes using ellipsoids that match the size distribution of our dataset. Unlike [21] and [22], our approach randomly samples healthy CT slices from an existing nodule-positive case rather than from an entirely novel healthy case. This strategy significantly reduces data dependency, particularly in data scarcity scenarios. For the downstream 2D lung nodule segmentation task, we randomly crop a $16\times 512\times 512$ patch from a nodule-positive CT image, ensuring that the selected slices do not contain lung nodules. We then generate a nodule mask within the automatically annotated lung regions. For the downstream 3D lung nodule segmentation task, we follow the same procedure as in the 2D task, but instead use a $128^{3}$ patch.
@@ 3.3.2 EAAS process
To integrate the proposed pulmonary DPM-solver with the existing anatomically aware sampling (AAS) process and further enhance model efficiency, we introduce the effective anatomically aware sampling (EAAS) method, which directly generates lung nodules. To address the imperfect blending issue between synthetic and background images, Lung-DDPM generates pulmonary regions containing lung nodules, which are then blended with extra-pulmonary areas. However, this approach is rigid and inefficient, as it only supports $128^{3}$ resolution due to computational constraints. To overcome this limitation, the EAAS process begins with random noise $n$, a reference CT patch $x$ cropped from a nodule-positive CT scan, and the corresponding synthetic nodule mask mm. The input size is $64^{3}$, allowing the EAAS process to be executed using a lighter model. An intermediate state $x_{t}$ is computed by mixing the inverted $x$ and the random noise $n$ under the constraint of $m$, enabling the denoising model to dynamically blend the nodule and its surrounding pulmonary region. This inversion process follows the basic forward diffusion procedure [12], which gradually adds Gaussian noise to the input $x$, generating a series of intermediate states $x_{1},\ldots,x_{t-1},x_{t}$ with a variance $\beta_{t}\in(0,1)$:
@@ 4.1.1 LIDC-IDRI
In our experiments, we firstly use the well-known pulmonary CT dataset LIDC-IDRI (Lung Image Database Consortium and Image Database Resource Initiative) [11], a widely used public benchmark for lung nodule detection, classification, and segmentation in thoracic CT scans. The dataset originally contains 1,018 anonymized low-dose chest CT scans with corresponding lung nodule annotations from four experienced radiologists, who marked and categorized nodules based on their characteristics and likelihood of malignancy.
To obtain unified nodule annotations, we apply a 50% consensus criterion [23] to manage the variability among the four radiologists. Since existing automatic segmentation methods are unstable for detecting lung nodules in contrast-enhanced CT images[24, 25, 26, 27], we exclude 509 contrast-enhanced cases from the original dataset. Additionally, we remove 25 cases containing multiple nodules and two cases with oversized nodules (>128>128mm).
@@ 4.1.2 NLST
To evaluate the generalization capability of our proposed approach, we utilized the publicly available National Lung Screening Trial (NLST) dataset [28], a large-scale, multi-center study comparing the effectiveness of low-dose computed tomography (LDCT) and chest X-ray (CXR) for lung cancer screening in high-risk populations. The NLST includes current or former smokers aged 55–74 years with a minimum 30 pack-year smoking history; former smokers were eligible if they had quit within the past 15 years. Notably, the original NLST dataset provides only binary malignancy labels, without explicit nodule annotations. To ensure experimental rigor and exclude benign cases with nodules, we leveraged nodule-positive cases identified by Cherezov et al. [29] and our previous work [30], extracting healthy slices from each case to generate synthetic nodules. In total, we selected 252 participants [^1] One case was excluded due to resampling failure. who underwent non-contrast CT scans. Nodule annotations were sourced from Cherezov et al. [29], with all nodules manually delineated by a radiologist. Note that the NLST dataset is only used for generating synthetic samples for the lung nodule segmentation evaluation.
@@ 4.2 Evaluation metrics
For the lung nodule segmentation evaluation, we perform 5-fold cross-validation with an 80% training / 20% test split. The proposed method and its competitors are trained using the training set from each fold and generate synthetic images based on synthetic nodule layouts derived from the training set. Finally, we train a U-Net [31] using a real-synthetic mixed training set and report the Dice score and Hausdorff distance (HD95) metrics, along with a 95% confidence interval computed across the five validation folds.
@@ 4.4.1 Efficiency evaluation
GEM-3D [9] 1.90 19.18 288.92 1.90 19.05 34.10
Med-DDPM [6] 4.35 6.15 53.33 2.02 6.13 45.89
Lung-DDPM [8] 4.35 6.17 52.93 2.02 6.16 45.42
@@
Table 2 presents the efficiency evaluation results, comparing Lung-DDPM+, Lung-DDPM [8], Med-DDPM [6], and GEM-3D [9] in 3D ($128^{3}$) and 2D ($16\times 512\times 512$) lung nodule synthesis tasks. The results clearly demonstrate that Lung-DDPM+ significantly outperforms all competitors in computational efficiency, achieving the lowest TFLOPs, VRAM usage, and sampling time. Specifically, Lung-DDPM+ requires only 0.54 TFLOPs, 0.91 GB of VRAM, and 3.79 seconds per sample for both 3D and 2D synthesis, which is up to 8× fewer TFLOPs, 21× less VRAM, and 76× faster sampling time compared to Lung-DDPM and Med-DDPM. In contrast, GEM-3D exhibits the highest computational cost, with 1.90 TFLOPs, 19+ GB of VRAM usage, and significantly longer sampling times (288.92s for 3D and 34.10s for 2D), making it the least efficient method. Med-DDPM and Lung-DDPM perform similarly, requiring 4.35 TFLOPs and over 6 GB of VRAM for 3D synthesis, while their sampling times (52.93s and 53.33s, respectively) remain far from the real-time feasibility achieved by Lung-DDPM+. These findings highlight that Lung-DDPM+ maintains competitive segmentation performance while offering vastly superior efficiency, making it particularly well-suited for resource-constrained and real-time medical imaging applications.

### v2

@@ 1 Introduction
Lung cancer is one of the greatest threats to human well-being. In 2022, it remained the most frequently diagnosed cancer worldwide, with approximately 2.5 million new cases, accounting for 12.4% of all cancer diagnoses. Furthermore, lung cancer continues to be the leading cause of cancer-related mortality, responsible for approximately 1.8 million deaths (18.7% of total cancer deaths) in 2022 [bray2024global]. For small, localized tumors (stage I), surgical resection of non-small cell lung cancer (NSCLC) offers a favorable prognosis, with 5-year survival rates of up to 70%. In contrast, cases diagnosed with distant metastatic disease (stage IV) have a 1-year survival rate of only 15-19%, compared to 81-85% for stage I [blandin2017progress]. These statistics underscore the critical importance of early detection of lung nodules and intervention in lung cancer diagnosis.
@@
Given the rapid advancement of AI in healthcare, numerous deep learning-based models have been deployed across various domains. However, this progress exacerbates data scarcity, which significantly impacts AI-assisted diagnostic models for early lung cancer detection. Recently, generative AI has garnered increasing attention for its ability to provide high-quality synthetic samples, enhancing model performance and mitigating data limitations. In radiological imaging, generative AI applications extend to X-ray [salehinejad2018synthesizing], computed tomography (CT) [jiang2020covid], positron emission tomography (PET) [luo2022adaptive], magnetic resonance imaging (MRI) [dorjsembe2024conditional], and ultrasound [liang2022sketch]. Among these, CT image synthesis remains one of the most challenging tasks due to its high computational complexity and the intricate texture details of anatomical structures.
@@
With the emergence of generative AI, various models for CT image synthesis have been introduced over the past decade. Since 2020, our team has focused on generative models for thoracic CT synthesis, initially proposing a generative adversarial network (GAN)-based approach for 2D COVID-19 CT image synthesis [jiang2020covid]. More recently, we advanced this work by introducing Lung-DDPM [jiang2025lungddpm], a denoising diffusion probabilistic model (DDPM) for 3D CT image synthesis with lung nodules. While Lung-DDPM addressed three key limitations of existing methods—anatomical imprecision, imperfect blending, and inconsistency—SOTA models still suffer from low efficiency in generating 3D thoracic CT images.
@@
As illustrated in Figure 1 (A), we compare the sampling times of the proposed Lung-DDPM+ with three SOTA models—Lung-DDPM [jiang2025lungddpm], Med-DDPM [dorjsembe2024conditional], and GEM-3D [zhu2024generative]—for generating thoracic CT images of two different sizes. To synthesize a single CT image, SOTA generative models require anywhere from tens to over two hundred seconds, resulting in low efficiency during the sampling process. This inefficiency severely limits the practical deployment of generative models in downstream medical applications, particularly in time-sensitive and cost-sensitive scenarios.
@@
Beyond the low efficiency issue, we also observed that SOTA generative models tend to synthesize imprecise anatomical structures when no semantic layout guidance is available. In Figure 1 (B), we illustrate a case generated by Lung-DDPM and Med-DDPM, both exhibiting anatomical imprecision. Compared to the proposed Lung-DDPM+ sample, the synthetic image produced by Med-DDPM failed to maintain anatomical precision in extra-pulmonary areas. While Lung-DDPM preserved extra-pulmonary regions, its synthetic sample lacked pulmonary vessel and bronchus details. In clinical practice, experienced radiologists manually annotate detailed semantic layouts for organs and lesions. However, this process is time-consuming and costly [ma2021toward], making well-labeled medical imaging datasets scarce for researchers. Therefore, developing a feasible approach to address the anatomical imprecision issue in scenarios with limited annotations remains a critical research challenge.
@@
    Our method has low data requirements and does not rely on externally sourced healthy cases, making it applicable to datasets in which all cases contain nodules (e.g., LIDC-IDRI [samuel2011lung]). By synthesizing additional nodule-positive samples with greater diversity, our approach enhances the robustness and performance of downstream segmentation models. While healthy cases are not required, they can be incorporated to further expand the diversity of the synthetic dataset.
@@ 2.1 Efficient denoising diffusion probabilistic models
Since the introduction of DDPM [ho2020denoising], it has emerged as one of the most advanced generative models due to its strong ability to generate high-quality data across various applications. However, low efficiency remains a critical limitation, restricting both its performance and deployment. To address this issue, researchers have explored multiple strategies to enhance DDPM efficiency. Improved model architectures, such as latent space diffusion models [rombach2022high], allow DDPMs to operate in a compressed space, significantly reducing computational demands. Adaptive noise scheduling methods, including PriorGrad [lee2021priorgrad], introduce data-dependent priors to accelerate convergence and improve sample quality. Additionally, hybrid model integration, which combines diffusion models with pre-trained generative architectures like GANs and VAEs [dhariwal2021diffusion], leverages the strengths of both approaches to enhance sampling speed. Further optimizations focus on the sampling process, where techniques such as efficient sampling schedules [watson2021learning] and early stopping mechanisms [lyu2022accelerating] reduce the number of denoising steps while maintaining generation quality. Despite these advancements, research on optimizing DDPM sampling efficiency in medical imaging remains limited, necessitating further exploration to improve its feasibility in clinical applications.
@@ 2.2 Generative models for medical imaging
Over the past decade, generative models have been extensively utilized in medical imaging, driven by rapid advancements in deep learning. For instance, Jiang et al. [jiang2020covid] employed a conditional generative adversarial network (cGAN) to synthesize COVID-19 CT images, facilitating data augmentation for improved diagnostic models. Dorjsembe et al. [dorjsembe2024conditional] introduced conditional diffusion models for semantic 3D brain MRI synthesis, enhancing the generation of realistic brain images for various clinical applications. Zhu et al. [zhu2024generative] proposed a generative enhancement approach for 3D medical images, focusing on improving image quality and resolution. Diffusion models have also gained traction in medical imaging, as highlighted in a recent review by Hein et al. [hein2024physics], which explores the transformative role of physics-inspired generative models in this domain. Among these advancements, Jiang et al. [jiang2025lungddpm] developed Lung-DDPM, a semantic layout-guided diffusion model for lung CT image synthesis, demonstrating superior performance over other state-of-the-art (SOTA) methods. However, like other approaches, it suffers from low efficiency, a key challenge this paper aims to address.
@@ 3.2 Pulmonary DPM-solver
Unlike Lung-DDPM, which operates on the entire pulmonary region, the proposed method focuses on the effective synthesis of nodule regions. To accelerate the anatomically aware sampling (AAS) process, we introduce a pulmonary DPM-solver, which extends the original DPM-solver [lu2022dpm, lu2022dpm++] by dynamically blending nodules with surrounding pulmonary areas during the EAAS sampling process.
@@
Despite their impressive generative capabilities, standard DPMs suffer from slow inference, often requiring hundreds or thousands of function evaluations (NFEs) for high-quality sampling. To mitigate this limitation, DPM-Solver [lu2022dpm] was introduced as a fast, high-order ODE solver that significantly accelerates sampling while preserving sample fidelity.
@@
Building upon DPM-Solver, DPM-Solver++ [lu2022dpm++] further optimizes inference speed and quality by integrating both stochastic and deterministic solvers in a hybrid framework. In early diffusion steps, stochastic updates enhance sample diversity, while deterministic high-order solvers refine results in later timesteps:
@@ 3.3.1 Nodule semantic layout generation
The generation of pulmonary CT images relies heavily on high-quality lung nodule semantic layouts, which are typically created by experienced radiologists in the Lung-DDPM workflow. To reduce radiologists’ workload and enhance the efficiency of the generation process, we introduce a rule-based nodule semantic layout generation method. Specifically, following [chen2024towards] and [hu2023label], we first generate realistic tumor-like shapes using ellipsoids that match the size distribution of our dataset. Unlike [chen2024towards] and [hu2023label], our approach randomly samples healthy CT slices from an existing nodule-positive case rather than from an entirely novel healthy case. This strategy significantly reduces data dependency, particularly in data scarcity scenarios. For the downstream 2D lung nodule segmentation task, we randomly crop a $16\times 512\times 512$ patch from a nodule-positive CT image, ensuring that the selected slices do not contain lung nodules. We then generate a nodule mask within the automatically annotated lung regions. For the downstream 3D lung nodule segmentation task, we follow the same procedure as in the 2D task, but instead use a $128^{3}$ patch.
@@ 3.3.2 EAAS process
To integrate the proposed pulmonary DPM-solver with the existing anatomically aware sampling (AAS) process and further enhance model efficiency, we introduce the effective anatomically aware sampling (EAAS) method, which directly generates lung nodules. To address the imperfect blending issue between synthetic and background images, Lung-DDPM generates pulmonary regions containing lung nodules, which are then blended with extra-pulmonary areas. However, this approach is rigid and inefficient, as it only supports $128^{3}$ resolution due to computational constraints. To overcome this limitation, the EAAS process begins with random noise $n$, a reference CT patch $x$ cropped from a nodule-positive CT scan, and the corresponding synthetic nodule mask mm. The input size is $64^{3}$, allowing the EAAS process to be executed using a lighter model. An intermediate state $x_{t}$ is computed by mixing the inverted $x$ and the random noise $n$ under the constraint of $m$, enabling the denoising model to dynamically blend the nodule and its surrounding pulmonary region. This inversion process follows the basic forward diffusion procedure [ho2020denoising], which gradually adds Gaussian noise to the input $x$, generating a series of intermediate states $x_{1},\ldots,x_{t-1},x_{t}$ with a variance $\beta_{t}\in(0,1)$:
@@ 4.1.1 LIDC-IDRI
In our experiments, we firstly use the well-known pulmonary CT dataset LIDC-IDRI (Lung Image Database Consortium and Image Database Resource Initiative) [samuel2011lung], a widely used public benchmark for lung nodule detection, classification, and segmentation in thoracic CT scans. The dataset originally contains 1,018 anonymized low-dose chest CT scans with corresponding lung nodule annotations from four experienced radiologists, who marked and categorized nodules based on their characteristics and likelihood of malignancy.
To obtain unified nodule annotations, we apply a 50% consensus criterion [kubota2011segmentation] to manage the variability among the four radiologists. Since existing automatic segmentation methods are unstable for detecting lung nodules in contrast-enhanced CT images[shaish2023diagnostic, christensen2021optimizing, ye2022deep, kihara2025impact], we exclude 509 contrast-enhanced cases from the original dataset. Additionally, we remove 25 cases containing multiple nodules and two cases with oversized nodules (>128>128mm).
@@ 4.1.2 NLST
To evaluate the generalization capability of our proposed approach, we utilized the publicly available National Lung Screening Trial (NLST) dataset [national2011national], a large-scale, multi-center study comparing the effectiveness of low-dose computed tomography (LDCT) and chest X-ray (CXR) for lung cancer screening in high-risk populations. The NLST includes current or former smokers aged 55–74 years with a minimum 30 pack-year smoking history; former smokers were eligible if they had quit within the past 15 years. Notably, the original NLST dataset provides only binary malignancy labels, without explicit nodule annotations. To ensure experimental rigor and exclude benign cases with nodules, we leveraged nodule-positive cases identified by Cherezov et al. [cherezov2018delta] and our previous work [jiang2025benchmark], extracting healthy slices from each case to generate synthetic nodules. In total, we selected 252 participants [^1] One case was excluded due to resampling failure. who underwent non-contrast CT scans. Nodule annotations were sourced from Cherezov et al. [cherezov2018delta], with all nodules manually delineated by a radiologist. Note that the NLST dataset is only used for generating synthetic samples for the lung nodule segmentation evaluation.
@@ 4.2 Evaluation metrics
For the lung nodule segmentation evaluation, we perform 5-fold cross-validation with an 80% training / 20% test split. The proposed method and its competitors are trained using the training set from each fold and generate synthetic images based on synthetic nodule layouts derived from the training set. Finally, we train a U-Net [ronneberger2015u] using a real-synthetic mixed training set and report the Dice score and Hausdorff distance (HD95) metrics, along with a 95% confidence interval computed across the five validation folds.
@@ 4.4.1 Efficiency evaluation
GEM-3D [zhu2024generative] 1.90 19.18 288.92 1.90 19.05 34.10
Med-DDPM [dorjsembe2024conditional] 4.35 6.15 53.33 2.02 6.13 45.89
Lung-DDPM [jiang2025lungddpm] 4.35 6.17 52.93 2.02 6.16 45.42
@@
Table 2 presents the efficiency evaluation results, comparing Lung-DDPM+, Lung-DDPM [jiang2025lungddpm], Med-DDPM [dorjsembe2024conditional], and GEM-3D [zhu2024generative] in 3D ($128^{3}$) and 2D ($16\times 512\times 512$) lung nodule synthesis tasks. The results clearly demonstrate that Lung-DDPM+ significantly outperforms all competitors in computational efficiency, achieving the lowest TFLOPs, VRAM usage, and sampling time. Specifically, Lung-DDPM+ requires only 0.54 TFLOPs, 0.91 GB of VRAM, and 3.79 seconds per sample for both 3D and 2D synthesis, which is up to 8× fewer TFLOPs, 21× less VRAM, and 76× faster sampling time compared to Lung-DDPM and Med-DDPM. In contrast, GEM-3D exhibits the highest computational cost, with 1.90 TFLOPs, 19+ GB of VRAM usage, and significantly longer sampling times (288.92s for 3D and 34.10s for 2D), making it the least efficient method. Med-DDPM and Lung-DDPM perform similarly, requiring 4.35 TFLOPs and over 6 GB of VRAM for 3D synthesis, while their sampling times (52.93s and 53.33s, respectively) remain far from the real-time feasibility achieved by Lung-DDPM+. These findings highlight that Lung-DDPM+ maintains competitive segmentation performance while offering vastly superior efficiency, making it particularly well-suited for resource-constrained and real-time medical imaging applications.